# Dynamical Lensing Tomography of Black Hole Ringdown

### v2

@@ A.2.1 QNM to Hertz potential
We calculate the quasinormal mode using the qnm package Stein (2019), which employs Leaver’s method to the radial equation and the Cook-Zalutskiy spectral approach to the angular sector Leaver (1986); Cook and Zalutskiy (2014). Furthermore, we use the pseudospectral method to get solutions to the radial Teukolsky equation. Here we briefly illustrate our steps. We convert the asymptotic behavior into coordinate $r$ Cook and Zalutskiy (2014)
$\lim_{r\rightarrow\infty}{}_{s}R_{\ell mn}(r)\sim r^{-1-2s+2i\omega_{mn}M}e^{i\omega_{mn}r}$ (58)
$\lim_{r\rightarrow r_{+}}{}_{s}R_{\ell mn}(r)\sim\left(r-r_{+}\right)^{-s-i\sigma_{+}}$ (59)
@@
${}_{s}R_{\ell mn}(r)=\left(r-r_{+}\right)^{-s-i\sigma_{+}}r^{-1-s+2i\omega_{mn}M+i\sigma_{+}}e^{i\omega_{mn}r}{}_{s}\tilde{R}_{\ell mn}(r)\,,$ (60)
@@
$\lim_{r\to\infty}{}_{s}\tilde{R}_{\ell mn}(r)=1$ (61)
We can expand ${}_{s}\tilde{R}_{\ell mn}(r)$ as a series of Chebyshev polynomials and get ${}_{s}S_{\ell mn}(\theta)$ in terms of spin-weighted spherical functions as in Ref. Cook and Zalutskiy (2014). As a result, we can use our numerical results and evaluate them for any point $(t,r,\theta,\phi)$ of spacetime in the subsequent ray tracing section. To ensure the perturbation metric is flawless, we also verified ^$\hat{\mathcal{E}}_{ab}(h)=0$ numerically.
A.3 Radiated energy
In this section, we estimate the radiated energy of a single mode after truncation. The energy flux is given by Berti et al. (2007)
@@ A.3 Radiated energy
where $t_{0}=r_{*}(r_{o})$. This total energy satisfies the condition $E_{\mathrm{tot}}/M\lessapprox 3\%$ Berti et al. (2007). Assuming that only the $\ell=m=2$ mode dominates the radiated energy, this relationship leads to
$|{}_{s}Z|\lessapprox 0.24\,.$ (64)
@@ Appendix B Backward Ray Tracing
B.1 Local oberver basis
@@
Within the Boyer-Lindquist coordinate system, a ZAMO frame can be chosen and expressed in terms of the coordinate basis $\left\{\partial_{t},\partial_{r},\partial_{\theta},\partial_{\phi}\right\}$ Cunha et al. (2016); Frolov and Novikov (1998); Johannsen (2013); Bardeen (1973) as
$\vec{z}_{(t)}^{\mu}=\frac{g_{\phi\phi}\partial_{t}-g_{\phi t}\partial_{\phi}}{\sqrt{g_{\phi\phi}\left(g_{\phi t}^{2}-g_{\phi\phi}g_{tt}\right)}}\,,\,\vec{z}_{(r)}^{\mu}=\frac{\partial_{r}}{\sqrt{g_{rr}}}\,,\,\vec{z}_{(\theta)}^{\mu}=\frac{\partial_{\theta}}{\sqrt{g_{\theta\theta}}}\,,\,\vec{z}_{(\phi)}^{\mu}=\frac{\partial_{\phi}}{\sqrt{g_{\phi\phi}}}\,.$ (65)
@@
$g^{(0)}_{\mu\nu}\vec{z}_{(\rho)}^{\mu}\vec{z}_{(\sigma)}^{\nu}=\eta_{(\rho)(\sigma)}\,,$ (66)
with $\eta_{(\rho)(\sigma)}=\operatorname{diag}\{-1,1,1,1\}$. Unlike the unperturbed case, Eq. (66) no longer holds when the unperturbed metric $g^{(0)}_{\mu\nu}$ is replaced by the perturbed metric $g_{\mu\nu}$. To address this issue, we employ a Gram-Schmidt orthogonalization process, using the chosen ZAMO tetrad $\vec{v}_{\rho}=\vec{z}_{(\rho)}^{\mu}$ as the initial set of vectors. The Gram-Schmidt process can be expressed as
$\vec{u}_{0}$ $=\vec{v}_{0}\,,$ (67)
@@
$\operatorname{proj}_{\vec{u}}\vec{v}=\frac{\langle\vec{u},\vec{v}\rangle}{\langle\vec{u},\vec{u}\rangle}\vec{u}\,,$ (68)
@@
$\tilde{\vec{z}}_{(\rho)}^{\mu}=\frac{\vec{u}_{\rho}}{\sqrt{\absolutevalue{\langle\vec{u}_{\rho}\,,\vec{u}_{\rho}\rangle}}}$ (69)
@@
$g_{\mu\nu}\tilde{\vec{z}}_{(\rho)}^{\mu}\tilde{\vec{z}}_{(\sigma)}^{\nu}=\eta_{(\rho)(\sigma)}\,.$ (70)
B.2 Impact Parameters
In the image plane of the observer defined above, each photon is assigned Cartesian coordinates $(\tilde{X},\tilde{Y})$, which represent its impact parameters. These coordinates are proportional to the photon’s observation angles $(\tilde{\alpha},\tilde{\beta})$. Here we follow a convention similar to Ref. Cunha et al. (2016), and define them as:
$\tilde{X}\equiv-r\tilde{\beta},\quad\tilde{Y}\equiv r\tilde{\alpha}\,.$ (71)
Considering the geometry of photon detection, we can express the photon’s 4-momentum in the observer’s reference frame as follows (refer to Fig. 1 in Ref. Cunha et al. (2016)):
$p^{(\phi)}=|\vec{P}|\sin\tilde{\beta}\cos\tilde{\alpha}\,,\,p^{(\theta)}=|\vec{P}|\sin\tilde{\alpha}\,,\,p^{(r)}=|\vec{P}|\cos\tilde{\beta}\cos\tilde{\alpha}\,,$ (72)
@@ B.2 Impact Parameters
$p^{\mu}=\tilde{\vec{z}}_{(\sigma)}^{\mu}p^{(\sigma)}\,.$ (73)
B.3 Kerr–Schild coordinates
To avoid the numerical problem when a photon gets too close to the pole in the BL coordinate system, we perform ray tracing in the Cartesian Kerr–Schild (KS) coordinate system $\{t_{\mathrm{KS}},x,y,z\}$. The coordinate transformations, which can be found in Ref. Christian and Chan (2021); Mädler and Winicour (2019); Font et al. (1999), are also reviewed in the following.
@@ B.3 Kerr–Schild coordinates
$r$ $=\left[\frac{R^{2}-a^{2}+\sqrt{\left(R^{2}-a^{2}\right)^{2}+4a^{2}z^{2}}}{2}\right]^{1/2},$ (76)
@@
$\phi_{\mathrm{KS}}=\phi+a\int\frac{dr}{\Delta_{\mathrm{BL}}}\,.$ (79)
Following Ref. Christian and Chan (2021), we fix the integration constant is fixed by letting $t_{\mathrm{KS}}=t$ and $\phi_{\mathrm{KS}}=\phi$ at the initial radius $r_{o}$, so we have
$t_{\mathrm{KS}}$ $=t+\frac{M}{\sqrt{M^{2}-a^{2}}}\left[r_{+}\ln\left(\frac{r-r_{+}}{r_{o}-r_{+}}\right)-r_{-}\ln\left(\frac{r-r_{-}}{r_{o}-r_{-}}\right)\right]\,,$ (80)

### v3

@@ 2.1 QNM to Hertz potential
We calculate the quasinormal mode using the qnm package [39], which employs Leaver’s method to the radial equation and the Cook-Zalutskiy spectral approach to the angular sector [95, 23]. Furthermore, we use the pseudospectral method to get solutions to the radial Teukolsky equation. Here we briefly illustrate our steps. We convert the asymptotic behavior into coordinate $r$ [23]
$\lim_{r\rightarrow\infty}{}_{s}R_{\ell mn}(r)\sim r^{-1-2s+2i\omega_{mn}M}e^{i\omega_{mn}r}$ (A.58)
$\lim_{r\rightarrow r_{+}}{}_{s}R_{\ell mn}(r)\sim\left(r-r_{+}\right)^{-s-i\sigma_{+}}$ (A.59)
@@
${}_{s}R_{\ell mn}(r)=\left(r-r_{+}\right)^{-s-i\sigma_{+}}r^{-1-s+2i\omega_{mn}M+i\sigma_{+}}e^{i\omega_{mn}r}{}_{s}\tilde{R}_{\ell mn}(r)\,,$ (A.60)
@@
$\lim_{r\to\infty}{}_{s}\tilde{R}_{\ell mn}(r)=1$ (A.61)
We can expand ${}_{s}\tilde{R}_{\ell mn}(r)$ as a series of Chebyshev polynomials and get ${}_{s}S_{\ell mn}(\theta)$ in terms of spin-weighted spherical functions as in Ref. [23]. As a result, we can use our numerical results and evaluate them for any point $(t,r,\theta,\phi)$ of spacetime in the subsequent ray tracing section. To ensure the perturbation metric is flawless, we also verified ^$\hat{\mathcal{E}}_{ab}(h)=0$ numerically.
3 Radiated energy
In this section, we estimate the radiated energy of a single mode after truncation. The energy flux is given by [37]
@@ 3 Radiated energy
where $t_{0}=r_{*}(r_{o})$. This total energy satisfies the condition $E_{\mathrm{tot}}/M\lessapprox 3\%$ [37]. Assuming that only the $\ell=m=2$ mode dominates the radiated energy, this relationship leads to
$|{}_{s}Z|\lessapprox 0.24\,.$ (A.64)
@@ Appendix B Backward Ray Tracing
1 Local oberver basis
@@
Within the Boyer-Lindquist coordinate system, a ZAMO frame can be chosen and expressed in terms of the coordinate basis $\left\{\partial_{t},\partial_{r},\partial_{\theta},\partial_{\phi}\right\}$ [33, 29, 83, 30] as
$\vec{z}_{(t)}^{\mu}=\frac{g_{\phi\phi}\partial_{t}-g_{\phi t}\partial_{\phi}}{\sqrt{g_{\phi\phi}\left(g_{\phi t}^{2}-g_{\phi\phi}g_{tt}\right)}}\,,\,\vec{z}_{(r)}^{\mu}=\frac{\partial_{r}}{\sqrt{g_{rr}}}\,,\,\vec{z}_{(\theta)}^{\mu}=\frac{\partial_{\theta}}{\sqrt{g_{\theta\theta}}}\,,\,\vec{z}_{(\phi)}^{\mu}=\frac{\partial_{\phi}}{\sqrt{g_{\phi\phi}}}\,.$ (B.65)
@@
$g^{(0)}_{\mu\nu}\vec{z}_{(\rho)}^{\mu}\vec{z}_{(\sigma)}^{\nu}=\eta_{(\rho)(\sigma)}\,,$ (B.66)
with $\eta_{(\rho)(\sigma)}=\operatorname{diag}\{-1,1,1,1\}$. Unlike the unperturbed case, Eq. (B.66) no longer holds when the unperturbed metric $g^{(0)}_{\mu\nu}$ is replaced by the perturbed metric $g_{\mu\nu}$. To address this issue, we employ a Gram-Schmidt orthogonalization process, using the chosen ZAMO tetrad $\vec{v}_{\rho}=\vec{z}_{(\rho)}^{\mu}$ as the initial set of vectors. The Gram-Schmidt process can be expressed as
$\vec{u}_{0}$ $=\vec{v}_{0}\,,$ (B.67)
@@
$\operatorname{proj}_{\vec{u}}\vec{v}=\frac{\langle\vec{u},\vec{v}\rangle}{\langle\vec{u},\vec{u}\rangle}\vec{u}\,,$ (B.68)
@@
$\tilde{\vec{z}}_{(\rho)}^{\mu}=\frac{\vec{u}_{\rho}}{\sqrt{\absolutevalue{\langle\vec{u}_{\rho}\,,\vec{u}_{\rho}\rangle}}}$ (B.69)
@@
$g_{\mu\nu}\tilde{\vec{z}}_{(\rho)}^{\mu}\tilde{\vec{z}}_{(\sigma)}^{\nu}=\eta_{(\rho)(\sigma)}\,.$ (B.70)
2 Impact Parameters
In the image plane of the observer defined above, each photon is assigned Cartesian coordinates $(\tilde{X},\tilde{Y})$, which represent its impact parameters. These coordinates are proportional to the photon’s observation angles $(\tilde{\alpha},\tilde{\beta})$. Here we follow a convention similar to Ref. [33], and define them as:
$\tilde{X}\equiv-r\tilde{\beta},\quad\tilde{Y}\equiv r\tilde{\alpha}\,.$ (B.71)
Considering the geometry of photon detection, we can express the photon’s 4-momentum in the observer’s reference frame as follows (refer to Fig. 1 in Ref. [33]):
$p^{(\phi)}=|\vec{P}|\sin\tilde{\beta}\cos\tilde{\alpha}\,,\,p^{(\theta)}=|\vec{P}|\sin\tilde{\alpha}\,,\,p^{(r)}=|\vec{P}|\cos\tilde{\beta}\cos\tilde{\alpha}\,,$ (B.72)
@@ 2 Impact Parameters
$p^{\mu}=\tilde{\vec{z}}_{(\sigma)}^{\mu}p^{(\sigma)}\,.$ (B.73)
3 Kerr–Schild coordinates
To avoid the numerical problem when a photon gets too close to the pole in the BL coordinate system, we perform ray tracing in the Cartesian Kerr–Schild (KS) coordinate system $\{t_{\mathrm{KS}},x,y,z\}$. The coordinate transformations, which can be found in Ref. [79, 86, 75], are also reviewed in the following.
@@ 3 Kerr–Schild coordinates
$r$ $=\left[\frac{R^{2}-a^{2}+\sqrt{\left(R^{2}-a^{2}\right)^{2}+4a^{2}z^{2}}}{2}\right]^{1/2},$ (B.76)
@@
$\phi_{\mathrm{KS}}=\phi+a\int\frac{dr}{\Delta_{\mathrm{BL}}}\,.$ (B.79)
Following Ref. [79], we fix the integration constant is fixed by letting $t_{\mathrm{KS}}=t$ and $\phi_{\mathrm{KS}}=\phi$ at the initial radius $r_{o}$, so we have
$t_{\mathrm{KS}}$ $=t+\frac{M}{\sqrt{M^{2}-a^{2}}}\left[r_{+}\ln\left(\frac{r-r_{+}}{r_{o}-r_{+}}\right)-r_{-}\ln\left(\frac{r-r_{-}}{r_{o}-r_{-}}\right)\right]\,,$ (B.80)
@@
Appendix C Results of Mirror Mode
In this section, we consider the effect of incorporating the QNM’s mirror mode with equal energy, reflecting a more realistic scenario where the SMBHB was in a quasi-circular orbit before the merger. In Fig. S2, we show the evolution of $\tilde{\delta}$ for different values of $\Delta\tilde{\rho}$ and $\theta_{o}$. Compared to the results of Figs. 4 and 5 in the maintext, this setup also exhibits the two previously discussed characteristics: exhibit the same two key features: an initial rapid decay during the ringdown phase, followed by a late-time power-law tail. Due to the mirror mode’s symmetry across the equatorial plane, the dependence on $\theta_{o}$ becomes less pronounced.
Figure S2: Evolution of the deflection angles of geodesics terminating at infinity, plotted as a function of observation time $\tilde{t}_{o}$, for the case incorporating the QNM’s mirror mode with equal energy. Left: Results for various impact parameters $\Delta\tilde{\rho}/M$ with $\theta_{o}/\pi=1-10^{-4}$. Right: Results for various initial inclination angles $\theta_{0}$ with $\tilde{\rho}/M=10$ along the $+\tilde{X}$ axis.